A BOD test is run using 100ml of treated wastewater mixed with 200 mL of pure water. The initial DO of the mix is 9 mg/L. After 5 days, the DO is 4 mg/L. What is the five day BOD of the wastewater?
1
Expert's answer
2020-12-14T14:43:04-0500

BOD = DO0 - DO1= 9-4=5 mg/L

Answer: 5 mg/L